Can you help me with the below?

3. Gains from trade Consider two neighboring island countries called Dolorium and Contente. They each have 4 million labor hours available per month that they can use to produce rye, jeans, or a combination of both. The following table shows the amount of rye or jeans that can be produced using 1 hour of labor. Rye Jeans Country (Bushels per hour of labor) (Pairs per hour of labor) Dolorium Contente 32 12 24 Initially, suppose Contente uses 1 million hours of labor per month to produce rye and 3 million hours per month to produce jeans, while Dolorium uses 3 million hours of labor per month to produce rye and 1 million hours per month to produce jeans. Consequently, Dolorium produces 24 million bushels of rye and 32 million pairs of jeans, and Contente produces 12 million bushels of rye and 72 million pairs of jeans. Assume there are no other countries willing to trade goods, so, in the absence of trade between these two countries, each country consumes the amount of rye and jeans it produces Doloriums opportunity cost of producing 1 bushel of rye is 4 pairs ▼ of eans, and Contentes opportunity cost of producing 1 bushel of rye is 2 pairs ▼ of jeans. Therefore, Contente ▼ has a comparative advantage in the production of rye, and Dolorum ▼ has a comparative advantage in the production of jeans Suppose that each country completely specializes in the production of the good in which it has a comparative advantage, producing only that good. In this case, the country that produces rye will produce illion bushels per month, and the country that produces jeans will produce million pairs per month. In the following table, enter each countrys production decision on the third row of the table (marked Production) Suppose the country that produces rye trades 26 million bushels of rye to the other country in exchange for 78 million pairs of jeans. In the following table, select the amount of each good that each country exports and imports in the boxes across the row marked Trade Action, and enter each countrys final consumption of each good on the line marked Consumption.
Answer #1

Dolorium's opportunity cost of producing 1 bushel of rye is 32/8 = 4 pairs of jeans and Contente's opportunity cost of producing 1 bushel of rye is 24/12 =2 pairs of jeans. Because Contente has a lower opportunity cost for the production of rye . Therefore, Contente has a comparative advantage in the production of rye and Dolorium has a comparative advantage in the production of jeans.

Suppose that each country completely specializes in the production of the good in which it has a comparative advantage , producing only that good . In this case,the country that produces rye i.e Contente will produce (4)(12) = 48 million bushels per month And the country that produces jeans i.e Dolorium will produce 4(32)=128 million pairs per month.

Suppose the country that produces rye trades 26 million bushels of rye to the other country in exchange for 78 million pairs of jeans.

DOLORIUM CONTENTE
Rye (milions of bushels) Jeans (Millions of pairs) Rye (Millions of bushels) Jeans (Millions of pairs)
Without trade Production 24 32 12 72
Consumption 24 32 12 72
With Trade Production 0 128 48 0
Trade action Import 26 Export 78 Export 26 Import 78
Consumption 26 (128-78)=50 (48-26)=22 78
Gains from trade Increase in consumption (26-24)= 2 (50-32)= 18 (22-12)=10 (78-72)= 6

When the two countries did not specialize ,the total production of rye was 36 million bushels per month and the total production of jeans was 104 millions pairs per month. Because of specialization ,the total production of rye was 48 million bushels per month and total production of jeans was 128 million pair per month. This implies that because of specialization , total production of rye has increased by (48-36)=12 million bushels per month and total production of jeans has increased by (128-104)=24 million pairs per month.